# BeetleVerse: A study on taxonomic classification of ground beetles

### v1

@@ 2 Related Work
The computational identification of insects has emerged as a critical research domain, with numerous studies addressing the challenges of taxonomic classification through advanced machine learning (ML) techniques [36, 49, 3, 51, 48, 5, 25]. Domain adaptation is critical [53, 43, 26, 15, 34]. For example, [43] investigated the generalization of ML models across different imaging conditions, revealing significant performance drops for out of distribution data. Several works proposed multimodal approaches integrating visual features with ecological metadata [11, 21, 2], showing different signal modalities often evolve independently to serve distinct functions. A few recent works have explored automated methods for identifying and monitoring carabids [6, 22, 18, 50, 1]. Most notable of them is NEON’s continental-scale ground beetle monitoring program [22, 37]. Among others, [6] achieved 84.6% species-level accuracy using linear discriminant analysis on pitfall trap-collected carabids; [18] achieved 51.9% species-level accuracy using applied neural networks; and [1] got 90% accuracy on Sri Lankan beetles by using background removal, segmentation, and transfer learning. However, these approaches are limited by small taxonomic coverage and the use of conventional ML models that lack the representational power of cutting-edge vision encoders. Our work addresses these limitations by leveraging that power and covering a broader taxonomic range.
@@ 3.1 Datasets used in this study
  • BeetlePUUM: A dataset of Hawaiian-endemic carabids, imaged from pinned specimens at the Pu’u Maka’ala site of National Ecological Observatory Network (NEON).
  • BeetlePalooza: Another dataset of NEON carabids, from 30 sites across the continental US, containing images of beetles preserved in ethanol vials [14].
@@ 3.3 Taxonomic Classification
Our primary focus is to evaluate the performance of the state-of-the-art models on taxonomic classification task for all the datasets. We conduct this by extracting feature embeddings from various pretrained foundation models and training a multi-layer perceptron (MLP) classifier on top of the embeddings. The embeddings come from diverse vision architectures categorized into: (1) vision-only self-supervised models (DINOv2 [39], ViTMAE [20], SwAV [16], MoCov3 [8]), (2) vision-only supervised models (SWINv2 [30], BeIT [4], LeViT [17], ConvNeXt[31]), and (3) vision+language supervised models (CLIP [40], BioCLIP [45], SigLIP [54], ViLT [27]). As a zero-shot baseline, we use BioCLIP[45], a CLIP-variant for fine-grained biological taxonomic classification.
@@ 4.2.1 Zero-Shot Prediction with BioClip
On first result evaluates zero-shot prediction of BioClip across three taxonomic levels (Table 3). Accuracy is highest at the family level and declines rapidly through the genus to the species level. This gradual drop is due to the model’s limited exposure to diverse taxonomic groups, as the train data for BioClip lacked representation of carabid taxa.
@@ 4.4 Cross-Dataset Domain Adaptation
Figure 4 illustrates a comparative overview of performance results of the top performing model ViLT across two adaptation scenarios, from which we see the that model generalizes well when trained on one lab dataset (NHM) and tested on another (BeetlePalooza), with a score of 0.91, but performs poorly when tested on a dataset with in-situ images (I1MC), with an average accuracy score of 0.57. The scores corroborate with the observation that cross-domain generalization in taxonomic classification is a persistent challenge [43]. We report the detailed performance scores in Table 12.
Figure 4: Cross-dataset domain adaptation performance of ViLT. Case A: Train on NHM, Test on I1M (at genus); Case B: same as A (at species); Case C: Train on BeetlePalooza, Test on I1M (at genus); Case D: same as C (at species); Case E: Train on NHM, Test on BeetlePalooza (at genus); Case X: Average of Case A, B, C, D. (All four are cases of Lab to In-situ domain shift)
The model’s poor performance on I1MC dataset refers to morphological and contextual variations between curated lab specimens and in-situ images, where uncontrolled variables like heterogeneous substrates (leaf litter, soil textures), dynamic lighting conditions, non-uniform specimen orientations (dorsal/lateral views), occlusions by debris or vegetation, moisture effects on cuticular reflectance, and presence of microorganisms (mites, fungi) collectively degrade model robustness across domains. This underscores the need for domain-adaptive strategies that can account for such visual noise and ecological variability.
@@ 5 Conclusion and Future Work
Our work presents a significant advancement in streamlining carabid taxonomy. However, there are critical gaps that need to be resolved. Firstly, although we compile data from four taxonomically distant sources, still it only covers $<$5% of all carabid species. Second, the domain-shift score drop on I1MC data reveals a major domain gap. In the future, we will work on bridging these gaps, as well as analyzing predicted taxa using phylogenetic and ecological proximity, and testing the models ability to flag a novel taxa. With these efforts, we hope to develop a scalable framework of taxonomic classification for carabids, advancing ecological analysis and biodiversity monitoring.
@@ Acknowledgment
This work is supported by both the Imageomics Institute and the AI and Biodiversity Change (ABC) Global Climate Center. The Imageomics Institute is funded by the US National Science Foundation’s Harnessing the Data Revolution (HDR) program under Award #2118240 (Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning). The ABC Global Climate Center is funded by the US National Science Foundation under Award No. #2330423 and Natural Sciences and Engineering Research Council of Canada under Award No. #585136. Additionally, this material is based upon work supported by the National Science Foundation under Award Numbers #2301322, #1950364, #242918, and #1926569, as well as the USDA National Institute of Food and Agriculture, Hatch Project Number ME0-22425 through the Maine Agricultural and Forest Experiment Station for SR and AE. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the US National Science Foundation, the US Department of Agriculture, or Natural Sciences and Engineering Research Council of Canada.
We also thank Tanya Berger-Wolf, Graham Taylor, Iuliia Eyriay, Scott Lowe, Lisa Wu, Mridul Khuarana, Mike Long, Evan Donoso, Elizabeth Campolongo, and Matthew Thompson for their contributions to BeetlePUUM and collegial discussions and support, which have enriched our perspectives for this work.
@@
Among the four datasets we used, all except BeetlePUUM are publicly available: BeetlePalooza [14], NHM-Carabids [19], I1MC [38].
@@ A.1 Dataset Details
This dataset digitize all pinned carabid specimens from the NEON Pu’u Maka’ala ecological observatory site in Hawaii. The collection was assembled from 600 original source images: 420 bulk images captured with a Canon EOS DSLR camera (model 7D with a 24-105 macro lens) and 180 high-detail microscopic images acquired using a SWIFCAM SC1603 system featuring a 16MP 1/2.33” CMOS sensor. The bulk images each contained 3-5 pinned specimens arranged vertically in sequential order according to their unique specimen IDs, with all individuals within a given image representing the same species and captured from the same pitfall trap. Then we run Grounding DINO [29] to precisely detect and crop out individual beetles from these group images. Each image is accompanied by comprehensive metadata including geolocation coordinates, collection date, and full taxonomic verification by carabid specialists. Using the coordinates and collection dates, we extract relevant weather data for each specimen, providing ecological context for morphological analyses. Morphological traits are measured using a digital annotation tool named TORAS [24]. Figure  5 displays a group image of carabid specimens and their corresponding individual crops.

### v2

@@ 2 Related Work
The computational identification of insects has emerged as a critical research domain, with numerous studies addressing the challenges of taxonomic classification through advanced machine learning (ML) techniques [39, 57, 3, 59, 55, 5, 28]. Domain adaptation is critical [61, 49, 29, 17, 37]. For example, [49] investigated the generalization of ML models across different imaging conditions, revealing significant performance drops for out of distribution data. Several works proposed multimodal approaches integrating visual features with ecological metadata [12, 24, 2], showing different signal modalities often evolve independently to serve distinct functions. A few recent works have explored automated methods for identifying and monitoring carabids [6, 25, 21, 58, 1, 45]. Most notable of them is NEON’s continental-scale ground beetle monitoring program [25, 40]. Among others, [6] achieved 84.6% species-level accuracy using linear discriminant analysis on pitfall trap-collected carabids; [21] achieved 51.9% species-level accuracy using applied neural networks; and [1] got 90% accuracy on Sri Lankan beetles by using background removal, segmentation, and transfer learning. However, these approaches are limited by small taxonomic coverage and the use of conventional ML models that lack the representational power of cutting-edge vision encoders. Our work addresses these limitations by leveraging that power and covering a broader taxonomic range.
@@ 3.1 Datasets used in this study
  • BeetlePUUM: A dataset of Hawaiian-endemic carabids, imaged from pinned specimens at the Pu’u Maka’ala site (PUUM) of National Ecological Observatory Network (NEON) [46]. It contains high quality images taken in controlled settings and is rich with ecological metadata, though limited in taxonomic diversity.
  • BeetlePalooza: Another dataset of NEON carabids, from 30 sites across the continental US, containing images of beetles preserved in ethanol vials [16]. Although it broadens taxonomic and geographic diversity (36 genera, 76 species) and contains ecological metadata, the specimens vary in spacing and orientation due to the collection’s scale and the fragile nature of the specimens.
@@ 3.3 Taxonomic Classification
Our primary focus is to evaluate the performance of state-of-the-art models on the taxonomic classification task across all datasets. To achieve this, we extract feature embeddings from various pretrained foundation models and train a multi-layer perceptron (MLP) classifier on top of these embeddings. The embeddings are derived from diverse vision architectures, categorized into: (1) vision-only self-supervised models (DINOv2 [42], ViTMAE [23], SwAV [19], MoCov3 [9]), (2) vision-only supervised models (SWINv2 [33], BeIT [4], LeViT [20], ConvNeXt [34]), and (3) vision+language supervised models (CLIP [43], BioCLIP [51], SigLIP [62], ViLT [30]). For example, the carabid Carabus vietinghoffii belongs to the family Carabidae, genus Carabus, and species vietinghoffii, illustrating the hierarchical structure used in our classification. As a zero-shot baseline, we use BioCLIP [51], a CLIP-variant for fine-grained biological taxonomic classification.
@@ 4.2.1 Zero-Shot Prediction with BioClip
Our first evaluation focuses on the zero-shot prediction performance of BioCLIP across three taxonomic levels (Table 3). Accuracy is highest at the family level and declines rapidly through the genus to the species level. This gradual drop is attributed to the model’s limited exposure to diverse taxonomic groups, as BioCLIP’s training data lacks sufficient representation of carabid taxa. Consequently, zero-shot results for I1MC and NHM-Carabids are omitted due to their limited taxonomic overlap with BioCLIP’s training data, which reduces its applicability to these diverse datasets and is partially reflected in the challenges observed with the merged dataset.
@@ 4.4 Cross-Dataset Domain Adaptation
Figure 4 illustrates a comparative overview of the performance of the top-performing model ViLT across two adaptation scenarios. The model generalizes well when trained on one lab dataset (NHM-Carabids) and tested on another (BeetlePalooza), achieving a score of 0.91 at the genus level, but performs poorly when tested on in-situ images (I1MC), with an average accuracy score of 0.57. This performance drop is driven by morphological and contextual variations, such as variable lighting, non-standard orientations (e.g., lateral views), heterogeneous substrates (e.g., leaf litter, soil textures), occlusions by debris or vegetation, moisture effects on cuticular reflectance, and the presence of microorganisms (e.g., mites, fungi) in I1MC, unlike the controlled dorsal views in NHM-Carabids. These scores highlight the persistent challenge of cross-domain generalization in taxonomic classification [49]. Detailed performance scores are reported in Table 12.
Figure 4: Cross-dataset domain adaptation performance of ViLT. Case A: Train on NHM-Carabids, Test on I1MC (at genus); Case B: same as A (at species); Case C: Train on BeetlePalooza, Test on I1MC (at genus); Case D: same as C (at species); Case E: Train on NHM-Carabids, Test on BeetlePalooza (at genus); Case X: Average of Cases A, B, C, D. Cases A to D represent lab-to-in-situ domain shifts for both genus and species levels, while Case E is evaluated at the genus level only due to limited species-level overlap between NHM-Carabids and BeetlePalooza.
The poor performance on I1MC underscores the need for domain-adaptive strategies to address ecological variability and visual noise, enhancing model robustness across diverse domains.
@@ 5 Conclusion and Future Work
Our work presents a significant advancement in streamlining carabid taxonomy. However, critical gaps remain to be addressed. Firstly, although we compile data from four taxonomically diverse sources, these still cover less than 5% of all carabid species. Secondly, the domain-shift score drop on I1MC data reveals a significant domain gap. While our study trains separate classifiers for genus and species levels, this approach may lead to inconsistent hierarchical predictions. To address the possibility of inconsistent hierarchical predictions for genus and species levels, we look forward to exploring hierarchical classification methods, such as Flamingo [8], or HAF [18], to ensure consistent taxonomic predictions by integrating the known taxonomy tree during training. Additionally, the limited taxonomic overlap between I1MC and other datasets (Table 2) highlights the potential for open-set recognition methods to address unseen classes, which we plan to investigate in future work. Moving forward, we aim to bridge these gaps by analyzing predicted taxa using phylogenetic and ecological proximity and testing the models’ ability to flag novel taxa. Through these efforts, we hope to develop a scalable framework for taxonomic classification of carabids, advancing ecological analysis and biodiversity monitoring.
@@ Acknowledgment
This work is supported by both the Imageomics Institute and the AI and Biodiversity Change Global Climate Center. The Imageomics Institute is funded by the US National Science Foundation’s Harnessing the Data Revolution program under Award #2118240 (Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning). The ABC Global Climate Center is funded by the US National Science Foundation under Award No. #2330423 and Natural Sciences and Engineering Research Council of Canada under Award No. #585136. Additionally, this material is based upon work supported by the National Science Foundation under Award Numbers #2301322, #1950364, #242918, and #1926569, as well as the USDA National Institute of Food and Agriculture, Hatch Project Number ME0-22425 through the Maine Agricultural and Forest Experiment Station for SR and AE.
@@
All four datasets are publicly available: BeetlePUUM [46], BeetlePalooza [16], NHM-Carabids [22], I1MC [41].
Code Availability Statement.
All codes to reproduce our results are available here.
@@ A.1 Dataset Details
This dataset digitize all pinned carabid specimens from the NEON Pu’u Maka’ala ecological observatory site in Hawaii. The collection was assembled from 600 original source images: 420 bulk images captured with a Canon EOS DSLR camera (model 7D with a 24-105 macro lens) and 180 high-detail microscopic images acquired using a SWIFCAM SC1603 system featuring a 16MP 1/2.33” CMOS sensor. The bulk images each contained 3-5 pinned specimens arranged vertically in sequential order according to their unique specimen IDs, with all individuals within a given image representing the same species and captured from the same pitfall trap. To ensure the images are optimized for advanced ecological applications like automated trait extraction, we follow the guidelines outlined in [13], which emphasize standardized specimen positioning, consistent size and color calibration, and comprehensive metadata documentation. Then we run Grounding DINO [32] to precisely detect and crop out individual beetles from these group images. The group images include comprehensive metadata comprising geolocation coordinates, collection dates, and taxonomic authentication by carabid specialists. Using the coordinates and date, we extract relevant weather data for each specimen, providing ecological context for morphological analyses. Morphological traits are measured using a digital annotation tool named TORAS [27]. Figure 5 displays a group image of carabid specimens and their corresponding individual crops.
@@
[47] presents a compilation of datasets BeetlePUUM and BeetlePalooza, featuring meticulously-measured morphological traits of individual specimens and offering a trait-based foundation for exploring taxonomic relationships and ecological variation among carabids. Beyond that, it serves as a valuable testbed for small-data regimes with multimodal data, a relatively less-explored area in ML where conventional approaches often underperform [56, 52].